OpenStudy (anonymous):

Find the right-handed limit: limit as x approaches pie/2 + e^tan(x)

OpenStudy (anonymous):

\(\infty\).

OpenStudy (anonymous):

I know infinity is a number, but my prof is stupied and shes like dont use it as a number, got into an arguement about it with her, so do i just say it does not exist?

OpenStudy (anonymous):

Your professor is right :D. \(\infty\) is NOT a number, but it gives an indication about how the expression grows as it approaches \(\pi/2\) in our case. So, yes you can also say the limit does not exist.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You're welcome.

myininaya (myininaya):

that is the dessert pie you have up there? can i have some?